OpenStudy (anonymous):

simplyfy it: ( 1+sinx)/cosx

Nnesha (nnesha):

do you know Pythagorean identity ??? if not its ok just let me know

OpenStudy (anonymous):

1/cos(x) + sin(x)/cos(x) csc(x) + tan(x)

OpenStudy (anonymous):

sec(x) + tan(x). My bad.

Nnesha (nnesha):

\[\sin = \frac{ o }{ h}\] ---\[\csc = \frac{ h }{ o}\] \[\cos = \frac{ a }{ h}\] ---\[\sec=\frac{ h }{ a}\] \[\tan = \frac{ o}{ a}\] \[\cot= \frac{ a}{ o}\]

Nnesha (nnesha):

so you can replace sin and cos by that  then solve